OpenStudy (anonymous):

Use the distributive property to factor the expression below. 54x^2y - 12x^2y^2

OpenStudy (anonymous):

oh easy what grade you in ? so i make sure i have the right formal

OpenStudy (anonymous):

Im not in a grade im in a program to graduate.

OpenStudy (anonymous):

give me a second then

OpenStudy (anonymous):

54x^2y = 2*3*3*3*x*x*y 12x^2y^2 = 2*2*3*x*x*y*y GCF = 2*3*x*x*y = 6x^2 so we get 6x^2y (9 - 2y) tada :P

OpenStudy (anonymous):

Thank you so much for showing me how to do it.

OpenStudy (anonymous):

no problem have a great day !